# SubTrack your Grad: Gradient Subspace Tracking for Memory and Time Efficient Full-Parameter LLM Training

### v1

@@ Appendix E Fine-Tuning RoBERTa
MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
Shared Parameters Batch Size 16 16 16 32 16 16 16 16
Learning Rate 1E-05 1E-05 3E-05 3E-05 1E-05 1E-05 1E-05 1E-05
@@
SubTrack-Grad SubTrack-Grad Step-Size 0.1 0.001 5.0 5.0 0.01 1.0 8.0 10.0
& GaLore, r = 4 Subspace Update Interval 500
Parameters $\alpha$ 4
SubTrack-Grad SubTrack-Grad Step-Size 0.1 0.1 5.0 13.0 0.1 1.0 5.0 10.0
& GaLore, r = 8 Subspace Update Interval 500
Parameters $\alpha$ 2
BAdam Parameters Block Switch Interval 100
@@
We also fine-tuned RoBERTa-Large on SuperGLUE tasks using the hyperparameters from Luo et al. (2024), as detailed in Table 13, with the exception that we fine-tuned each task for 30 epochs.
@@
SubTrack-Grad & SubTrack-Grad Step-Size 0.1 1.0 50.0 50.0 1 1.0
GaLore Parameters Subspace Update Interval 500 100 100 500 250 100
@@
Appendix F Pre-Training LLama-Based Architectures
To manage resources, we pre-trained all five Llama-based architectures for 10,000 iterations using hyperparameters reported in Table 14. While larger models typically require more iterations, this setup was sufficient for comparing methods rather than creating perfectly optimized models.
@@ Appendix F Pre-Training LLama-Based Architectures
60M 130M 350M 1B 3B
Architectural Hidden 512 768 1024 2048 4096
Parameters Intermediate 1376 2048 2736 5461 11008
Heads 8 12 16 24 32
Layers 8 12 24 32 32
Shared Parameters Learning Rate 1e-3 1e-3 1e-3 1e-5 1e-5
Batch Size 128 128 64 8 8
@@
SubTrack & Rank 128 256 256 512 512
GaLore Parameters Subspace Update Interval 200
SubTrack-Grad Step-Size 10
@@
Appendix G Time and Performance Consistency
Table 15 demonstrates that increasing the update frequency significantly increases GaLore’s runtime, while the performance of both methods remains comparable. This underscores SubTrack-Grad’s efficiency in reducing runtime without sacrificing performance. Notably, this advantage becomes even more pronounced for tasks that demand more frequent updates.
Table 15: Comparing changes in performance while decreasing the subspace update interval (increasing the frequency of subspace updates) on fine-tuning RoBERTa-Base on COLA task.
500 450 400 350 300 250 200 150 100 50
GaLore’s Wall-Time 458.79 460.19 471.36 483.04 495.83 520.28 556.14 605.1 715.87 1038.79
GaLore’s Performance 0.5358 0.5285 0.5385 0.5356 0.5309 0.54821 0.5332 0.5387 0.5463 0.5463
SubTrack-Grad’s Wall-Time 408.22 408.44 408.47 409.25 409.17 409.76 410.29 411.77 413.35 417.07
SubTrack-Grad’s Performance 0.5332 0.5382 0.5364 0.5385 0.5354 0.5364 0.5416 0.5385 0.5364 0.5673
Appendix H Convergence vs. Wall-Time
Figure 5 depicts the changes in the training loss function relative to wall-time. The results demonstrate that SubTrack-Grad’s wall-time reduction has minimal impact on the learning process, showcasing a convergence pattern comparable to other methods without introducing significant computational overhead.
Figure 5: The figures present changes in training loss relative to wall-time for pre-training the Llama-3B architecture on the C4 dataset and fine-tuning RoBERTa-Large on the BoolQ dataset. They demonstrate that SubTrack maintains the overall learning process for both pre-training and fine-tuning while avoiding significant computational overhead.

### v2

@@ Appendix E Fine-Tuning Experiments
SST-2 MRPC CoLA RTE STS-B
Shared Parameters Batch Size 16 16 32 16 16
@@
Low-Rank Optimizer Learning Rate 2E-05 2E-05 1E-05 2E-05 3E-05
Methods Parameters SubTrack Step-Size 0.01 3.0 5.0 15.0 10.0
Subspace Update Interval 500
Rank Config 8
$\alpha$ 2
BAdam Parameters Learning Rate 2E-05 2E-05 1E-05 2E-05 3E-05
Block Switch Interval 100
@@
Low-Rank Optimizer SubTrack++ Step-Size 0.1 1.0 50.0 100.0 1.0 1.0
Methods Parameters Subspace Update Interval 500 100 100 500 250 100
@@
The hyperparameters for fine-tuning RoBERTa-base are detailed in Table 6, matching those reported in the GaLore [Zhao et al., 2024] for rank-$8$ subspaces, with a subspace update interval set at 500 iterations. We also fine-tuned RoBERTa-Large on SuperGLUE tasks using the hyperparameters from Luo et al. [2024], as detailed in Table 7, with the exception that we fine-tuned each task for 30 epochs.
Appendix F Memory and Time Comparison
Table 8 presents the the peak memory consumption measured to compare SubTrack++ and other baselines. It shows that SubTrack++ requires on-par or better memory compared to GaLore [Zhao et al., 2024] and Fira [Chen et al., 2025] while getting state-of-the-art results. As detailed in Table 2, all baselines except BAdam [Luo et al., 2024] use the same number of optimizer parameters; therefore, any differences in their peak memory consumption stem from variations in their runtime parameters.
Table 8: Peak memory consumption comparison when pre-training Llama-based architectures with different sizes on C4 dataset. The 7B models are trained using the 8-bit Adam optimizer, except for the runs marked with . SubTrack++ demonstrates better or on-par memory compared to other low-rank methods that allows full-parameter training.
60M 130M 350M 1B 3B 7B
r=128 r=256 r=256 r=512 r=512 r=1024
Full-Rank 16.86 25.32 28.67 18.83 34.92 50.50
BAdam [Luo et al., 2024] 13.34 20.01 16.45 9.18 14.75 22.35
GaLore [Zhao et al., 2024] 16.89 25.52 27.85 14.74 26.03 36.00
Online Subspace Descent Liang et al. [2024] 16.61 25.86 28.76 18.45 32.57 33.10
LDAdam [Robert et al., 2025] 17.18 25.94 28.03 18.45 32.57 OOM
Fira [Chen et al., 2025] 16.39 24.99 27.33 14.68 25.62 47.84
SubTrack++ (Ours) 16.40 25.06 27.42 15.22 25.54 49.82
Additionally, Table 9 presents the wall-time consumed by each baseline across all model architectures during pre-training. Each run is configured to include exactly 10 subspace updates for SubTrack++ and other baselines employing periodic subspace updates. Specifically, models ranging from 60M to 3B use a subspace update interval of 200, resulting in 2,000 total iterations, while the 7B models use an interval of 500, yielding 5,000 iterations. Experiments for the 60M to 3B models are conducted on an NVIDIA A100 GPU, while the 7B model experiments are run on an NVIDIA RTX A6000.
Table 9: Wall-time comparison for pre-training LLaMA-based architectures of varying sizes on the C4 dataset. The number of iterations is set to ensure 10 subspace updates for methods using periodic subspace updates: 2k for models ranging from 60M to 3B, and 5k for the 7B model. SubTrack++ achieves the lowest wall-time among all baselines that support full-parameter training on large models. The 7B models are trained using the 8-bit Adam optimizer, except for the runs marked with . Since this can impact wall-time comparisons, it is more appropriate to compare runs within the same cluster.
60M 130M 350M 1B 3B 7B
r=128 r=256 r=256 r=512 r=512 r=1024
Full-Rank 524.0 1035.1 1396.4 974.9 1055.9 12726.2
BAdam [Luo et al., 2024] 511.3 779.2 961.6 798.6 1004.1 7283.7
GaLore [Zhao et al., 2024] 547.8 1094.2 1589.0 1729.5 2715.5 21590.4
Online Subspace Descent [Liang et al., 2024] 662.8 1228.2 1818.6 1438.7 1676.9 21590.4
LDAdam [Robert et al., 2025] 639.9 1342.2 2083.4 2780.9 4625.4 OOM
Fira [Chen et al., 2025] 635.3 1180.8 1729.7 1938.5 2898.4 22554.3
SubTrack++ (Ours) 627.6 1140.6 1593.2 1304.3 1517.6 16491.7
Appendix G Pre-Training LLama-Based Architectures
We pre-trained all six Llama-based architectures for 10,000 iterations using hyperparameters reported in Table 10. Hyperparameters are selected based on reported values in GaLore paper [Zhao et al., 2024].
@@ Appendix G Pre-Training LLama-Based Architectures
60M 130M 350M 1B 3B 7B
Architectural Hidden 512 768 1024 2048 2560 4096
Parameters Intermediate 1376 2048 2736 5461 6848 11008
Heads 8 12 16 24 32 32
Layers 8 12 24 32 32 32
Shared Parameters Learning Rate 1e-3 1e-3 1e-3 1e-4 1e-4 1e-4
Batch Size 128 128 64 8 8 4
Gradient Accumulation 2 2 2 2 2 4
@@
Low-Rank Optimizer Rank 128 256 256 512 512 1024
Methods Parameters Subspace Update Interval 200 200 200 200 200 500
SubTrack++ Step-Size 10
@@
Appendix H Subspace Tracking vs. SVD
Replacing SVD with Grassmannian-based subspace tracking, as opposed to the approach used in GaLore [Zhao et al., 2024] and Fira [Chen et al., 2025], offers two key advantages: (1) Robustness: As shown in Figure 5, Grassmannian tracking better handles noise and abrupt shifts, preserving subspace quality under dynamic conditions. (2) Efficiency: It significantly reduces computational overhead (see Table 2 and Appendix D). Figure 6 illustrates the wall-time of the experiments in the ablation in Figure 3, highlighting substantial speedups.
As a result, SubTrack++ achieves SOTA-level performance with GaLore-like memory use, while being one of the fastest methods; clearly demonstrated in Figure LABEL:fig:modular and 6.
Figure 6: Ablation study comparing pure Grassmannian subspace tracking with incremental additions of the projection-aware optimizer and recovery scaling, leading to SubTrack++ . Grassmannian tracking exceeds GaLore’s convergence, and significantly reduces wall-time, as also shown in Figure LABEL:fig:modular.